Dynamic Systems and Applications 15 (2006) 1-2

JOHN V. BAXLEY AND HIS CAREER AS A TEACHING MATHEMATICIAN

JOHN R. GRAEF^{1} AND JOHNNY HENDERSON^{2}

1Department of Mathematics, University of Tennessee at Chattanooga Chattanooga, TN 37403 USA. E-mail: John-Graef@utc.edu

2Department of Mathematics, Baylor University Waco, TX 76798 USA. E-mail: Johnny Henderson@baylor.edu

John V. Baxley received his Ph.D. from the University of Wisconsin under the direction of Seymour Parter in 1966, and to Parter’s chagrin, decided not to pursue a career at Stanford or Cornell, and choose instead to accept a position at Duke. He taught there for two years and decided to move to a then relatively unknown small North Carolina school, Wake Forest, where he remained for thirty-six years until his retirement in the spring of 2004.

John’s love of teaching and of mathematical research lead him down a path on which he refers to himself as a hybrid, namely a “teaching mathematician.” His research interests span a broad range of areas including Toeplitz operators, numerical solution of differential equations, operators in a Hilbert space, modeling of reactions in a chemostat, mathematical economics, and the area in which he is probably best known, namely, boundary value problems for nonlinear differential equations. Mathematical economics holds a special place in John’s work, and in fact, together with John Moorhouse, a colleague in the Economics Department at Wake Forest and Carroll Professor of Economics there, he created the Mathematical Economics major in 1975. This was one of the first such programs in the country.

But it is his love and dedication to teaching and mentoring students that drives John even more than his interest in doing research. And it is this of which he is the most proud. You do not have to be around John very long to feel this. In the Spring of 2004, just a few days before a retirement dinner for John was to take place at Wake, John Graef visited Baxley and gave a colloquium lecture. After the talk, Baxley excitedly came to the front of the room, not because lecture and results were that exciting, because they certainly were not, but because in his talk Graef made

Received April 15, 2005                 1056-2176 $15.00 \begin{matrix}c \\ ◯ \end{matrix}Dynamic Publishers, Inc.

2 JOHN R. GRAEF AND JOHNNY HENDERSON

reference to a paper that Baxley wrote with one of his students, Lyndsey J. Haywood. The fact that references were made to some of Baxley’s other papers in the lecture did not matter.

Over the years, John has co-authored more than a dozen research papers with his master’s thesis students and even with some of his undergraduate students. When discussing his relationship with his students, John once said

“We establish a mathematical bond which is very real and leads to a strong affection. When my students leave, they take a part of me with them, but they also leave a part of themselves with me. It is this tie that means the most to me.”

We dedicate this issue of Dynamic Systems and Applications to John on the occasion of his retirement after thirty-six years at Wake Forest University. The authors of these papers include some of the leading, and somewhat older, mathematicians in the field today as well as some younger researchers who will carry on the work in the years to come. John wouldn’t have it any other way.